OpenStudy (anonymous):

can someone find the derivative of (x^2)cos(1/x)+e^2

OpenStudy (anonymous):

youre gonna need the chain rule and product rule for the first part.\[\frac{d}{dx} (x^{2}\cos(\frac{1}{x})) = 2xcos(\frac{1}{x})+x^{2}(-\sin(\frac{1}{x}))(-\frac{1}{x^{2}})\] Then you gotta simplify

OpenStudy (anonymous):

cos(1/x)(2x)+(x^2)(sin(1/x)(x^-2) and the e cancles cause its a constant